OpenStudy (anonymous):

pyramid-shaped gum with a square base, the company has decided to place their pyramid gum inside a clear glass giant sphere. Each piece of gum has a base measurement of 1 inch and a height of 0.75 inches. The glass sphere container has a diameter of 17.25 inches. How many pieces of Pharaoh Chewing Gum can fit inside the glass sphere? You must show all work to receive credit. gum: b=1 h=0.75 v=1/3(1)(0.75) v=0.75/5/3= 0.25 glass sphere d=17.25 v=4/3(3.14)(8.625^3) v=268.24 now we divide the gum from the sphere 2.686/0.25=10.74 pieces of gum would fit. what am i doing wrong?

OpenStudy (anonymous):

my teacher says i should be using this but i dint get that answer when i use that formula The volume of the sphere would be V = (3.14)r^3 V = (3.14)(8.625^3) = 2686.25 in

OpenStudy (anonymous):

please i really need help